OpenStudy (anonymous):

if x≠0 , then u/x + 5u/x - u/5x= A. 7x/5u B. 5u/7x C. 29u/5x D. 31u/5x

OpenStudy (pawanyadav):

Take 5x LCM Then solve

OpenStudy (pawanyadav):

So I think C is correct

OpenStudy (anonymous):

whats the lcm of 5x?

OpenStudy (pawanyadav):

LCM of x,x,5x is 5x (all denominator)

OpenStudy (anonymous):

The terms may not be in school room order. \[-\frac{u x}{5}+\frac{5 u}{x}+\frac{u}{x}=\frac{30 u-u x^2}{5 x} \]